HCI + Na2SO4 > NaCI + H2SO4

A. What is my theoretical yield of NaCI if I start with 300 g of Na2SO4?

B. If my percent yield was 67%, what was my actual yield?
1
Expert's answer
2020-02-27T09:05:22-0500

The given reaction is mentioned below

"HCl + Na_2SO_4 \longrightarrow NaCl + H_2SO_4"

we will write balanced chemical equation for it

"2HCl + Na_2SO_4 \longrightarrow 2NaCl + H_2SO_4"

here , one mole of "Na_2SO_4" will give 2 mole of NaCl

gram molecular mass of Na2SO4 is equal to = 142 g/mole

gram molecular mass of NaCl = 58.5 g/mole

A.

here , given mass of Na2SO4 is 300 g

so mole of "Na_2SO_4 = \frac{given mass}{molar mass} = \frac {300}{142} = 2.113"

as per above balanced equation , 1 mole of Na2SO4 give 2 mole of HCl

so 2.113 mole of Na2SO4 will give "= 2\times 2.113" mole of NaCl

so, mass of NaCl produced="2\times2.113\times58.5=247.22 g" of NaCl

This is the theoretical yield in above case

B.

As it is given that percent yield was 67 %

so let x be the actual yield

now as per question

"67" % of x = 247.22 g

so , x"=247.22 \times \frac{100}{67}" = 368.98 g

so,actual theoretical yield was = 368.98 g of NaCl